If a^2+b^2+c^2= 20, a+b+c= 0 then find the value of ab+bc+ca?

Grade:9

2 Answers

Chelsi Kothari
6 years ago
(a+b+c)= a2 + b2+c2+2(ab + bc+ca)
let  (ab + bc+ca) be x
(0)= 20 + 2(x)
-20 =2x
-20/2 = x
-10 = x
Therefore ab+ bc+ca = -10
ankit singh
3 years ago
 
(a+b+c)= a2 + b2+c2+2(ab + bc+ca)
let  (ab + bc+ca) be x
(0)= 20 + 2(x)
-20 =2x
-20/2 = x
-10 = x
Therefore ab+ bc+ca = -10